OpenStudy (boxman61):

need couple a work checks if no one minds... Simplify: (4y+7)^{2} = 16y^{2}+49 x^{2}+4x-2(5x^{2}-x+3) = -9x^{2}+6x-6 7-3(x+4)=8(x-1) --> x=3/11

OpenStudy (boxman61):

(4y+7)^2=42y^2+72=16y^2+49 x^2+4x−2(5x^2−x+3)=x^2+4x−10x^2+2x−6=−9x^2+6x−6 solve: 7-3(x+4)=8(x-1) --> 7-3x-12=8x-8 --> -3x-5=8x-8 --> -11x=-3 --> x=3/11 that my calculations, just need a confirmation if possible, thank you one second ago Type your reply

OpenStudy (asnaseer):

\[(4y+7)^2=(4y+7)(4y+7)=16y^2+56y+49\]

OpenStudy (boxman61):

thank you

OpenStudy (asnaseer):

yw